OpenStudy (anonymous):

A right triangle has a hypotenuse of 13 cm and one leg that measures 12 cm. What is the length of the other leg?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

the second most common right triangle with integer sides, right behind 3-4-5 if 5-12 - 13

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

that is how i know without any computation. but with computation you use pythagoras and write \[12^2+b^2=13^2\] \[144+b^2=169\] \[b^2=25\] \[b=5\]

OpenStudy (anonymous):

or with a little practice you write \[b=\sqrt{13^2-12^2}\] but the computation is the same

OpenStudy (anonymous):

:)